### v2

# Linz, Austria

Abstract
We present f4ncgb, a new open-source C++ library for Gröbner basis computations in free algebras, which transfers recent advancements in commutative Gröbner basis software to the noncommutative setting. As our experiments show, f4ncgb establishes a new state of the art for noncommutative Gröbner basis computations. We also discuss implementation details and design choices.
1 Introduction
Over the last decades, Gröbner bases have become an indispensable tool in computational algebra and related fields. Originally introduced by Bruno Buchberger in his seminal PhD thesis [9], Gröbner bases offer an algorithmic framework for solving a wide range of problems in polynomial algebra. While initially developed for commutative polynomials, the theory has since been extended to various noncommutative contexts, including free (associative) algebras [4, 35]. In this setting, Gröbner bases (sometimes also referred to as Gröbner-Shirshov bases [7]) have found important applications in (linear) control theory [19], automated theorem proving [20, 10, 37, 21], as well as in graph [32] and game theory [41].
All these applications crucially rely on the ability to efficiently compute Gröbner bases in free algebras. While software for commutative Gröbner basis computations has seen remarkable progress in recent years, driven by the development of specialized data structures and optimized algorithms, noncommutative tools seem to lag behind. They often lack the same level of efficiency and sophistication, and mostly rely on outdated algorithms; see Section 2 for a more detailed discussion.
In this work, we present f4ncgb, a new open-source C++ library for Gröbner basis computations in free algebras, distributed under the license MIT and available at https://gitlab.sai.jku.at/f4ncgb/f4ncgb. The library supports noncommutative polynomial systems with coefficients in prime fields of characteristic $2^{31}$ and in the field of rational numbers. To the best of our knowledge, f4ncgb is the first open-source tool to port the newest advancements in Gröbner basis computations from the commutative setting to the noncommutative world. Our experiments in Section 8 show that f4ncgb establishes a new state of the art for noncommutative Gröbner basis computations among freely available tools.
The foundation of our implementation are specialized data structures for representing and storing noncommutative monomials and polynomials. We abstract these data types into a store structure, using a contiguous memory area for storing elements and a hash map for efficient lookup of known entries. We describe this structure in more detail in Section 5.
Algorithmically, the core of f4ncgb is an efficient implementation of Faugère’s F4 algorithm. Originally developed for commutative polynomials [14] and later adapted to free algebras [28, 40], the F4 algorithm reduces Gröbner basis computations to normal form computations of large sparse matrices via Gaussian elimination.
To efficiently support this, we have implemented a specialized Gaussian elimination algorithm tailored for sparse matrices, based on ideas from the commutative case [34]. Our implementation makes use of modern multi-core hardware by parallelizing matrix reductions, resulting in significant overall speedups. For inputs with rational coefficients, we employ a multi-modular strategy: computations are carried out modulo different primes, followed by a reconstruction step to recover the rational solution. We provide an overview of the noncommutative F4 algorithm in Section 4 and discuss our linear algebra implementation in Section 6.
In addition to computing Gröbner bases, f4ncgb offers proof logging capabilities in the form of cofactor representations: for each polynomial in the output basis, a representation in terms of the input can be computed, see Section 7. These representations are not only of interest for verifying the correctness of the output, but are also useful in applications [23].
2 Related Software
There already exists a wide range of tools for Gröbner basis computations in free algebras. In this section, we provide an overview of related software packages documented in the literature.
While the landscape of tools is diverse, almost all available packages share a common problem: they lag behind in terms of algorithmic development. In the commutative setting, essentially all modern Gröbner basis implementations rely on Faugère’s F4 algorithm, which provides substantial performance improvements over earlier methods. By contrast, almost all noncommutative packages continue to use the older and generally less efficient Buchberger algorithm.
Notable examples of tools relying on Buchberger’s algorithm include the Mathematica library NCAlgebra [18], the ApCoCoA package NCPoly [29], and the GAP package GBNP [11]. Other tools that also fall into this category are the Standard Lisp implementation bergman [3] and the C++ tool Opal [17], which both seem no longer actively maintained.
An important exception is the subscription-based computer algebra system Magma [8], which is one of the few tools to implement a noncommutative version of the F4 algorithm. In addition, Magma also features dedicated data structures and highly optimized subroutines (e.g., for the linear algebra), making it arguably the state of the art for computing noncommutative Gröbner bases. However, a significant drawback is its proprietary nature, which limits its availability to the community.
For homogeneous ideals over prime fields, the open-source system Macaulay2 [16] offers a decent implementation of the F4 algorithm. For all other ideals, however, it falls back to Buchberger’s algorithm.
To the best of our knowledge, the only open-source implementation of the noncommutative F4 algorithm over the field of rational numbers is the SageMath package operator_gb [5], developed by one of the authors. While operator_gb already implements the noncommutative F4 algorithm, it still relies on relatively simple data structures for representing and manipulating noncommutative objects. Additionally, the package is primarily written in interpreted Python, with selected components (such as the linear algebra) in compiled Cython, which limits performance. As a result, it cannot match the efficiency of the highly optimized and fully compiled C code in Magma. The new package f4ncgb can be viewed as a significant advancement of operator_gb, offering both improved algorithmic design and a more efficient implementation.
Special mention should also be given to the Letterplace [33] subsystem of Singular, which has established itself as the standard for freely available noncommutative Gröbner basis software. It offers a good tradeoff between efficiency and functionality, supporting a range of methods that go well beyond Gröbner basis computation alone (see [33] for an overview). Letterplace is based on the letterplace correspondence [30], a technique that embeds noncommutative computations into a larger commutative polynomial ring. This embedding allows for the reuse of optimized commutative data structures and subroutines in the noncommutative setting.
We see two notable limitations in the letterplace approach. The first one is inherent to the letterplace correspondence itself: the number of variables in the commutative ring grows linearly with the degree of the considered noncommutative polynomials. As a result, for high-degree inputs, the commutative computation can become prohibitively large. The second limitation is an implementation issue of Singular, as it still relies on Buchberger’s algorithm even for commutative computations.
3 Preliminaries – Noncommutative Gröbner Bases
To keep this work self-contained, we recall the most important aspects of the Gröbner basis theory in free algebras. For further details including proofs, we refer interested readers to, e.g., [40, 36, 21].
We write $\langle X\rangle$ for the free monoid on a finite set of indeterminates $X=\{x_{1},\ldots,x_{n}\}$. Recall that $\langle X\rangle$ consists of all finite words over the alphabet $X$, including the empty word $1$. Multiplication in $\langle X\rangle$ is given by concatenation of words, which also induces a natural notion of divisibility: for $v,w\in\langle X\rangle$, we say that $v$ divides $w$ if $v$ is a subword of $w$, i.e., if there exist $a,b\in\langle X\rangle$ such that $w=avb$.
For a field $K$, we consider the free (associative) algebra $K\langle X\rangle$. Elements in $K\langle X\rangle$ are finite sums of the form $c_{1}m_{1}+\cdots+c_{d}m_{d}$ with nonzero $c_{1},\dots,c_{d}\in K$ and pairwise different $m_{1},\dots,m_{d}\in\langle X\rangle$. We consider these elements as (noncommutative) polynomials with coefficients in $K$ and monomials in $\langle X\rangle$.
For a set of polynomials $F\subseteq K\langle X\rangle$, we let $(F)$ be the (two-sided) ideal generated by $F$, which consists of all two-sided linear combinations of the elements in $F$ with polynomials as coefficients, i.e., $(F)=\left\{\sum_{i=1}^{k}p_{i}f_{i}q_{i}\mid f_{i}\in F,\;p_{i},q_{i}\in K\langle X\rangle,\;k\in\mathbb{N}\right\}$. A set $F$ is a generating set of an ideal $I\subseteq K\langle X\rangle$ if $I=(F)$. Recall that if $|X|>1$, then there exist ideals in $K\langle X\rangle$ that do not possess a finite generating set. A classical example is the ideal $(\{xy^{n}x\mid n\in\mathbb{N}\})\subseteq K\langle x,y\rangle$.
A total ordering $\prec$ on $\left\langle X\right\rangle$ is called a monomial (well-)ordering if the following conditions hold:
  1. 1.
    $\forall a,b,v,w\in\left\langle X\right\rangle$: if $v\prec w$, then $avb\prec awb$;
  2. 2.
    every nonempty subset of $\langle X\rangle$ has a least element.
A very general class of monomial orderings are weighted block orderings of the following form: Fix $d\in\mathbb{N}_{>0}$ and assign to each variable $x_{i}\in X$ a nonzero weight vector $w_{i}\in\mathbb{R}_{\geq 0}^{d}$. Define the weight $w(m)$ of a monomial $m=x_{i_{1}}\dots x_{i_{k}}\in\langle X\rangle$ as $w(m)=\sum_{j=1}^{k}w_{i_{j}}\in\mathbb{R}_{\geq 0}^{d}$. The weighted block ordering $\prec_{\omega}$ w.r.t. the weight matrix $\omega=(w_{1},\dots,w_{n})$ is given by:
$m\prec_{\omega}m^{\prime}\quad\text{iff}\quad w(m)<w(m^{\prime})\text{ or }\left(w(m)=w(m^{\prime})\text{ and }m\prec_{\textup{lex}}m^{\prime}\right),$
where the real-valued vectors $w(m)$ and $w(m^{\prime})$ are compared lexicographically and $m\prec_{\textup{lex}}m^{\prime}$ denotes a lexicographic comparison of the words $m$ and $m^{\prime}$ from left to right.
At the moment, our implementation supports weighted block orderings where each weight vector is a unit vector $(0,\dots,0,1,0,\dots,0)$. This is sufficient to represent classical orderings such as degree-lexicographic orderings. It also enables the use of elimination orderings [21, Def. 2.4.15], which are crucial, e.g., for computing elimination ideals [21, Thm. 2.4.43].
For a fixed monomial ordering, the leading monomial $\operatorname{lm}(f)$ and leading coefficient $\operatorname{lc}(f)$ of a nonzero polynomial $f\in K\langle X\rangle$ are defined in the usual way. In our definition, the leading monomial does not contain a coefficient.
Definition 1
A Gröbner basis of an ideal $I\subseteq K\langle X\rangle$ (w.r.t. a monomial ordering $\prec$) is a subset $G\subseteq I$ such that for every nonzero $f\in I$, there exists $g\in G$ such that $\operatorname{lm}(g)$ divides $\operatorname{lm}(f)$.
We note that all standard characterizations of Gröbner bases from the commutative setting carry over naturally to the noncommutative case; see, e.g., [21, Thm. 2.4.37]. In particular, Buchberger’s criterion [13, Thm. 2.6.6], which provides an effective test for whether a set of commutative polynomials forms a Gröbner basis, has a noncommutative analogue in Bergman’s diamond lemma [4, Thm. 1.2].
To state this result, we recall the concept of ambiguities as introduced in [4]. In the following, we use $\lvert w\rvert$ to denote the length of a word $w\in\langle X\rangle$.
Definition 2
Let $f,g\in K\langle X\rangle\setminus\{0\}$ and let $a,b,c,d\in\langle X\rangle$ be such that one of the following conditions holds:
$\textit{1.}~\operatorname{lm}(af)$ $=\,\operatorname{lm}(gd)$ $\text{ and $b=c=1$ and $1\leq\lvert a\rvert<\lvert\operatorname{lm}(g)\rvert$}$,
$\textit{2.}~\operatorname{lm}(fb)$ $=\,\operatorname{lm}(cg)$ $\text{ and $a=d=1$ and $1\leq\lvert b\rvert<\lvert\operatorname{lm}(g)\rvert$}$,
$\textit{3.}~\operatorname{lm}(afb)$ $=\,\operatorname{lm}(g)$ $\text{ and $c=d=1$ and $f\neq g$},$
$\textit{4.}~\operatorname{lm}(f)$ $=\,\operatorname{lm}(cgd)\;$ $\text{ and $a=b=1$ and $f\neq g$}.$
Then we call the tuple $(a\otimes b,c\otimes d,f,g)$ an ambiguity of $f$ and $g$.
Note that two elements can give rise to multiple ambiguities, but only to finitely many. A polynomial may also form ambiguities (of type 1 and 2) with itself. If $\operatorname{\mathfrak{a}}=(a\otimes b,c\otimes d,f,g)$ is an ambiguity, then $\operatorname{lm}(afb)=\operatorname{lm}(cgd)$ by construction. Based on this, we define the degree of $\operatorname{\mathfrak{a}}$ to be $\deg(\operatorname{\mathfrak{a}})=\lvert\operatorname{lm}(afb)\rvert=\lvert\operatorname{lm}(cgd)\rvert$.
Analogous to Buchberger’s criterion, Bergman’s diamond lemma is based on a notion of (noncommutative) polynomial reduction, which is (basically) identical to the reduction used in the commutative case: $f$ can be reduced by $g$ if $\operatorname{lm}(g)$ divides a monomial in $f$ and the result of the reduction is $f-\lambda vgw$ with $\lambda\in K$, $v,w\in\langle X\rangle$ chosen so that the targeted monomial cancels. For a formal definition, see, e.g., [21, Sec. 2.4.2].
An ambiguity of two polynomials $f$ and $g$ reflects the fact that the monomial $\operatorname{lm}(afb)=\operatorname{lm}(cgd)$ can be reduced in two (possibly different) ways by $f$ and $g$. A set $G$ containing $f$ and $g$ can only be a Gröbner basis if these two reductions lead to the same canonical normal form. To test whether this is the case, one considers the associated S-polynomial.
Definition 3
Let $G\subseteq K\langle X\rangle$ and let $\operatorname{\mathfrak{a}}=(a\otimes b,c\otimes d,f,g)$ be an ambiguity of $f,g\in G$. The S-polynomial of $\operatorname{\mathfrak{a}}$ is
$\operatorname{\textup{S-Pol}}(\operatorname{\mathfrak{a}})=\frac{1}{\operatorname{lc}(f)}afb-\frac{1}{\operatorname{lc}(g)}cgd.$
Theorem 3.1(Bergman’s diamond lemma)
A generating set $G$ of an ideal $I\subseteq K\langle X\rangle$ is a Gröbner basis of $I$ if and only if, for all ambiguities $\operatorname{\mathfrak{a}}$ of $G$, the S-polynomial $\operatorname{\textup{S-Pol}}(\operatorname{\mathfrak{a}})$ reduces to 0 by $G$.
Theorem 3.1 provides an effective criterion to test whether a given set forms a Gröbner basis. It also directly gives rise to a completion procedure for constructing a Gröbner basis from an arbitrary generating set – a noncommutative analogue of Buchberger’s algorithm; see [21, Alg. 1] for a detailed description.
However, in contrast to the commutative case, not all finitely generated ideals in the free algebra admit a finite Gröbner basis (w.r.t. a fixed monomial ordering). For instance, the principal ideal $(xyx-xy)\subseteq\mathbb{Q}\langle x,y\rangle$ does not admit a finite Gröbner basis for any choice of ordering.
As a consequence, we cannot expect such a completion algorithm to terminate in general. Instead, we have to content ourselves with an enumeration procedure with the following behaviour:
  1. 1.
    If the input ideal $I$ admits a finite Gröbner basis (w.r.t. the chosen monomial order), the procedure terminates and returns such a basis.
  2. 2.
    Otherwise, it enumerates an infinite sequence $(g_{n})_{n\in\mathbb{N}}$ such that the set $\{g_{n}\mid n\in\mathbb{N}\}$ forms a Gröbner basis of $I$.
4 Noncommutative F4 Algorithm
Our library f4ncgb implements the noncommutative F4 algorithm for computing Gröbner bases in free algebras. Below, we provide a brief overview of the algorithm; for a detailed exposition, see [40, Sec. 5.4]. For those familiar with the commutative algorithm, we note that the noncommutative version follows the same high-level structure.
After initializing the prospective Gröbner basis with the input polynomials and computing all ambiguities between them, the F4 algorithm enters a loop consisting of the following main steps:
  1. 1.
    Critical pair selection: Select a subset of ambiguities, usually all those of minimal degree. For each selected ambiguity $\operatorname{\mathfrak{a}}=(a\otimes b,c\otimes d,f,g)$, form the polynomials afbafb and cgdcgd.
  2. 2.
    Symbolic preprocessing: Iterate over all monomials appearing in the polynomials generated in the previous step. For each monomial $m$, check whether it is divisible by the leading monomial of a basis element $g$. If so, form a polynomial vgwvgw with $v,w\in\langle X\rangle$ such that $\operatorname{lm}(vgw)=m$. If multiple choices for $v,g,w$ are possible, pick one (for $g$, e.g., the one with minimal leading monomial). Repeat this process also for all monomials in the newly created polynomials.
  3. 3.
    Matrix construction: Collect all polynomials constructed in the previous two steps in a matrix. To this end, fix an arbitrary order of the polynomials $f_{1},f_{2},\dots$ and order all monomials appearing in those polynomials in descending order $m_{1}\succ m_{2}\succ\dots$. The matrix entry at position $(i,j)$ is then given by the coefficient of monomial $m_{j}$ in polynomial $f_{i}$.
  4. 4.
    Matrix reduction: Reduce the matrix constructed in the previous step to (reduced) [^1] In principle, a row echelon form suffices. However, typically using the reduced row echelon form improves the algorithm’s overall performance. row echelon form.
  5. 5.
    Basis update: Translate the rows of the reduced matrix back into polynomials and add all those polynomials to the basis whose leading monomial is not divisible by an element of the current basis. Then, form all new ambiguities. Optionally, one can apply the noncommutative variant of the Gebauer-Möller criteria [40, Thm. 4.2.22] to discard redundant ambiguities. If ambiguities remain, continue with the next iteration.
This procedure enumerates a Gröbner basis in the sense of the previous section. To guarantee termination in practice, we have to add additional constraints. One way of doing this is by simply stopping the computation after a prescribed number of iterations. Another way is to impose an upper bound on the degree of ambiguities that are considered. If, during the execution of the algorithm, an ambiguity arises whose degree is larger than the designated bound, this ambiguity is discarded. Both approaches ensure termination but only yield a partial Gröbner basis. Our implementation supports both termination options.
5 Basic Data Structures – Monomials and Polynomials
Our library is based on specialized data structures for handling and storing noncommutative monomials and polynomials. The core operations that these data structures have to support are:
  1. 1.
    For monomials: comparison w.r.t. monomial ordering, multiplication, and divisibility tests.
  2. 2.
    For polynomials: leading monomial extraction and multiplication by monomials.
In our implementation, we represent each variable by a positive integer, which are assigned in increasing order according to the given monomial ordering, i.e., the smallest variable is assigned $1$, the next smallest $2$, and so on. Each noncommutative monomial then corresponds to a unique sequence of integers, with the empty sequence denoting the empty monomial $1$. This encoding allows lexicographic comparison between monomials to reduce to a fast lexicographic comparison of integer sequences.
To store monomials (and also polynomials), we have designed a common base store structure that is extended into monomial store and polynomial store structures using the Curiously Recurring Template Pattern [12]. The store implements a bump-pointer allocator combined with an open-addressing hash map (using Boost’s unordered_flat_map) for efficient lookups of known entries.
Each element in a store consists of some associated metadata followed by the data of the element itself. The metadata contains the length of the element (for a polynomial this is the number of terms) and optionally additional data.
Both the metadata and the stored values of each element are correctly aligned according to the requirements of their contained primitive data types, positionally allocated and deallocated, and automatically padded. We visualize the store’s structure in Figure 1.
[…]
Figure 1: Illustration of a monomial and polynomial store with global coefficient array over $\mathbb{Q}\langle x,y,z\rangle$, using a monomial ordering with $x\prec y\prec z$. The monomial store contains $x$, yxyx, xyxxyx, and xyzxyz (in this order). The polynomial store contains the polynomial $\frac{1}{2}xyz+2xyx-yx+\frac{2}{3}x$, referencing monomials by index. The indices above the monomial and polynomial store are used throughout our program to reference the objects.
We rely on malloc to return uninitialized memory in order to allocate $4\text{\,}\mathrm{GB}$ of sequential heap space when a store is initialized. This uninitialized memory is not yet hardwired into our process, instead its memory range is only reserved in our process space, still respecting memory limits. New entries are appended at the tip of the store, filling the allocated heap, initializing the memory, and actually wiring it into our process pagewise. This approach ensures that each appended element has a constant address, making pointers to it stable. We can therefore use pointers to elements during the entire program runtime.
For a Gröbner basis computation, we maintain one global monomial store that contains all monomials generated during the computation. Each monomial is uniquely identified by its index in this store, allowing for efficient referencing. When a new monomial is encountered, we first query the hash map to check whether it already exists in the store. If it does, we retrieve its index; if not, the sequence of integers representing the monomial is inserted into the store, along with its metadata, and assigned a new index.
[…]
Our store data structure allows for efficient storing and referencing of monomials (and polynomials), but it does not support fast divisibility tests – an operation frequently required in the F4 algorithm. In particular, monomial divisibility tests occur during symbolic preprocessing (step 2) and when computing ambiguities (step 5). In both cases, we have to test whether a monomial is divisible by one of the leading monomials of the current Gröbner basis.
To handle this efficiently, we use a separate data structure: a prefix tree (or trie). In such a tree, each node represents a variable, and each path from the root to a leaf encodes a monomial. When inserting a monomial, we traverse or construct the corresponding path based on its sequence of variables. Since common prefixes are shared, a prefix tree naturally supports fast prefix-based queries such as divisibility tests.
We maintain a global prefix tree consisting of all leading monomials of the current Gröbner basis, which is used for all divisibility tests. As elements are never removed from the basis – and hence from the prefix tree – the tree can be stored in contiguous memory, improving cache locality and enabling faster traversals.
5.1 Further Optimizations
Scratch Space for Products
In order for new elements to be looked up in a store’s hash map, they have to be constructed. Instead of using a separate buffer and a copy operation for this, we introduce a scratch space at the tip of a store’s heap. Each time a product of multiple monomials is computed, the resulting monomial is constructed in that space. After constructing the element, the store checks whether the element is already contained using its hash map. If it is, the scratch is discarded and later overwritten. If it is not, we commit the scratch space, storing the newly created element. We also update the hash map accordingly, increase the element count, and move the tip of the store to after the newly inserted element, including potential padding.
Memory Allocator and Tunables
[…]
5.2 Detrimental Optimizations
Saving Known Monomial Products
Instead of the scratch space, we also tried to cache the results of monomial multiplications in a product hash map. However, this approach required more memory and resulted in a slowdown compared to using the scratch space alone. While still surprising, this result can be explained by the reduced memory usage and improved cache locality of the scratch space method.
Compact Store Membership
To check whether an element already exists in the store, we use the store’s hash map. We tried to replace that hash map by a specialized hasher. Instead of hashing the elements directly, we store only the ID of each element from the store in a set. To test whether a new element in the scratch space shall be committed to the store, we then check for set membership in this set, testing not for equality of the IDs, but instead of the data they point to. This, however, dramatically increased the time for store membership tests in our experiments.
6 Linear Algebra
Computationally, the core operation of the F4 algorithm lies in the reduction of huge, structured sparse matrices to (reduced) row echelon form (step 1). This step dominates the overall runtime and thus demands a highly optimized implementation. In f4ncgb, we have implemented a specialized and parallelized Gaussian elimination procedure that exploits the structure of the matrices arising in Gröbner basis computations.
For computations over prime fields, our implementation closely follows the algorithm introduced in [34]. In the case of rational coefficients, we adopt a multi-modular approach, reducing rational computations to computations modulo different primes followed by a reconstruction step.
Since matrices arising in Gröbner basis computations typically only contain a few nonzero entries per row, we store them in a sparse row-wise format as an array of pointers to rows. Each row is itself composed of two arrays: one containing the column indices of the nonzero entries (sorted in ascending order), and the other containing the corresponding values.
6.1 Gaussian Elimination in Prime Fields
Our implementation supports prime fields with characteristic $2^{31}$. This allows us to represent matrix entries using 32-bit unsigned integers, which ensures both memory efficiency and fast arithmetic operations.
To compute the reduced row echelon form of an $m\times n$ matrix $A$ over $\mathbb{Z}_{p}$ for a prime $p<2^{31}$, we proceed in two steps: First, we transform $A$ into row echelon form, and then apply backward elimination to obtain the reduced row echelon form.
The forward elimination phase begins with a preprocessing step in which we permute the rows of $A$ such that, in each row, the first nonzero entry appears directly below or to the left of the first nonzero entry in the row above. This reordering yields a matrix with an approximately triangular structure, as illustrated in Figure 2. We note that this preprocessing step is crucial for the efficiency of the algorithm.
currently processed row(dense buffer)already knownpivot rows (sparse)
Figure 2: Visualization of our Gaussian elimination over $\mathbb{Z}_{p}$. Potentially nonzero entries are shaded in grey.
We then process the reordered matrix row by row starting from the top. To make row reductions more efficient, we maintain an auxiliary array pivots of length $n$ to keep track of already identified pivot elements. In particular, pivots[j] stores the index of the row containing a pivot in column $j$, or 1-1 if no pivot has been assigned to that column yet. This allows constant-time retrieval of pivot rows for any given column.
When processing a row $r$, we first identify the column index $j$ of its first nonzero entry. If column $j$ has no existing pivot, i.e, $\texttt{pivots[j]}=-1$, we do not perform any reductions on $r$. Instead, we update pivots[j] to reflect that row $r$ provides a pivot for column $j$, we normalize $r$ so that its pivot is $1$, and then proceed to the next row. Otherwise, we fully reduce row $r$ using all previously identified pivots.
As also done in commutative F4 implementations [34, 6], we perform row reductions using a dense data structure. In particular, we copy the current sparse row $r$ into a dense buffer – an auxiliary array of length $n$, initialized to 0. We then iterate over all positions of this buffer in ascending order, reducing all nonzero entries for which a pivot is known.
The buffer stores entries as 64-bit (signed) integers, which allows us to delay modular reductions and accumulate multiple arithmetic operations before applying expensive modulo operations. To prevent overflows during this accumulation, we add a square of the field characteristic $p$ if a value becomes negative. This is efficiently implemented via the branch-free update buffer[j] += (buffer[j] >> 63) & p*p.
Once the buffer is fully reduced, we copy its content back into the sparse row $r$. If the resulting row is nonzero, we normalize it to make the pivot equal to $1$ and update the pivots array accordingly.
Remark 1
[…]
Once all rows have been processed, the matrix is in row echelon form. To complete the computation of the reduced row echelon form, we first permute the rows again to restore the staircase shape illustrated in Figure 2, which may have been disrupted during the forward elimination phase. In this step, all zero rows are moved to the bottom. Then we process the rows again from top to bottom, now reducing each row completely by all rows above it. For this, we use the same dense reduction strategy as in the forward phase.
To take advantage of modern multi-core hardware, our implementation supports parallel Gaussian elimination, following the approach described in [34]. Whenever a row is ready to be reduced, this reduction is launched as a separate task and added to a global work queue, which is processed in parallel by a thread pool. Each thread is given its own local dense buffer to work in and all threads share access to a global pivots array. After reducing a row, if the result is nonzero, we use an atomic compare-and-swap (CAS) operation (realized via C++’s atomic_compare_exchange_weak) to update the global pivots array. If this CAS operation fails, meaning that another thread has already registered a pivot at the same position, we reload the row into the dense buffer and reduce it further.
Mersenne Primes
[…]
v1 = v + 1;
z = ((v1 >> b) + v1) >> b;
y = (v + z) & p;
Table 1 summarizes the benefits of this optimization on different CPUs and for different compilers. In our benchmark setting used in Section 8 (AMD EPYC 7313 with GCC 13.3.0), it yields a speedup of $3.5\times$. We have also tried to use [1, Alg. 5] for modular reductions by pseudo-Mersenne primes of the form $2^{b}-c$ with $c>1$ but this did not improve performance over the standard modulus operator.
CPU Compiler Standard [s] Optimized [s] Speedup
Intel i7-10700 GCC 15.1.1 11.685 66711.685\,667 1.311 9821.311\,982 8.906 880 68.906\,880\,6
Intel i7-10700 Clang 20.1.3 11.623 47511.623\,475 1.375 2721.375\,272 8.451 7678.451\,767
Intel Xeon w5-2455X GCC 13.3.0 4.709 9254.709\,925 1.177 1091.177\,109 4.001 265 04.001\,265\,0
AMD EPYC 7313 Clang 18.1.3 4.057 5754.057\,575 1.159 5511.159\,551 3.499 263 93.499\,263\,9
AMD EPYC 7313 GCC 13.3.0 4.057 7344.057\,734 1.159 7541.159\,754 3.498 788 53.498\,788\,5
Intel Xeon w5-2455X Clang 18.1.3 4.708 2474.708\,247 3.460 9933.460\,993 1.360 374 61.360\,374\,6
MacBook Air M2 (ARM64) AppleClang 16 1.263 3101.263\,310 1.491 2931.491\,293 0.847 123 940.847\,123\,94
Table 1: Speedup of Mersenne prime optimization on different CPUs when reducing all integers in the range $2^{31}\pm 2^{16}$ modulo $p=2^{31}-1$.
6.2 Multi-Modular Approach over the Rationals
Gröbner basis computations over the rational numbers typically suffer from severe intermediate coefficient growth, which significantly impacts performance. To address this issue, multi-modular techniques have been developed [2, 27]. These methods reduce computations over $\mathbb{Q}$ to computations modulo different primes, thereby bounding the size of the coefficients, and subsequently reconstruct the rational results from their modular images.
In the commutative setting, multi-modular methods are typically used in an end-to-end fashion: a complete Gröbner basis is computed modulo different primes and only the final rational Gröbner basis is reconstructed. It has been shown [22] that this classical strategy fails in the noncommutative case because of the infinite nature of Gröbner bases in free algebras.
Nevertheless, while an end-to-end multi-modular computation is not viable in our setting, the linear algebra within the F4 algorithm can still be performed in a multi-modular fashion. Our implementation uses the multi-modular algorithm presented in [38, Alg. 7.6] for computing the reduced row echelon form of rational matrices efficiently. In this way, we can still circumvent expensive rational arithmetic and leverage fast modular computations. Moreover, we highlight that end-to-end methods lack a general and efficient test to verify the correctness of the reconstructed result and therefore yield correct results only with high probability [2, 6]. In contrast, our method provides a simple verification criterion (see line 10 of Alg. 1) and is thus guaranteed to return the correct result.
The multi-modular algorithm is summarized in Algorithm 1. In this description, we denote by $H(A)=\max_{i,j}\lvert A_{i,j}\rvert$ the height of an integer matrix $A$. For a correctness proof of Algorithm 1, we refer to [38, Alg. 7.6]. For further details on the reconstruction via Chinese remaindering and rational reconstruction, see, e.g., [15, Ch. 5].
Algorithm 1 Multi-modular Reduced Echelon Form
0: A matrix $A\in\mathbb{Q}^{m\times n}$.
0: The reduced row echelon form of $A$.
[…]
2: choose a finite set $P$ of primes
3:while True do
4:  for $p\in P$ do
5:   $R_{p}\leftarrow$ the reduced row echelon form of $A^{\prime}\text{ mod }p$
6:  discard any $R_{p}$ whose pivot column vector is not maximal  [^2] The pivot column vector of $R_{p}$ consists of the indices of all pivot columns of $R_{p}$ sorted in ascending order. The maximality of such a vector is determined as follows: first the dimensions are compared (shorter vectors are smaller) and vectors of the same dimension are ordered according to the reverse lexicographic order.among all $R_{p}$, $p\in P$
7:  try to reconstruct from the modular echelon forms $R_{p}$, $p\in P$, a rational matrix $R$ via Chinese remaindering and rational reconstruction
8:  if the reconstruction succeeds then
9:   $d\leftarrow$ the smallest positive integer such that dRdR has only integer entries
10:   if $H(dR)\cdot H(A^{\prime})\cdot n<\prod_{p\in P}p$ then
11:    return $R$
12:  enlarge $P$ with primes not used so far
Tracer
To optimize Algorithm 1, we have implemented a technique described in [6]: During the very first modular computation in line 5, we construct a data structure called tracer, which records all rows of the matrix that reduce to zero. In all subsequent modular echelon form computations, we then apply this tracer to immediately zero out the corresponding rows, avoiding any additional computations. This optimization drastically speeds up the multi-modular algorithm, however the final output will only be correct with high probability. In particular, the result is guaranteed to be correct if the true reduced echelon form does not contain an entry whose denominator is divisible by the prime used to compute the tracer. This happens with high probability and the number of bad primes is finite. In our implementation, we use the Mersenne prime $p=2^{31}-1$ as the first prime for which the tracer is computed. This allows us to exploit the fast Mersenne modular operation discussed at the end of Section 6.1. The tracer can be disabled to ensure correctness of the output.
7 Proof Logging
Providing (efficient) certificates that a given set of polynomials forms a Gröbner basis remains an open problem. Aside from verifying the diamond lemma (Thm. 3.1) – which essentially amounts to re-running the Gröbner basis algorithm – no general method is currently known. We can, however, at least certify that each output polynomial computed by our implementation belongs to the ideal generated by the input using cofactor representations.
Definition 4
A cofactor representation of a polynomial $g\in K\langle X\rangle$ with respect to a set $F\subseteq K\langle X\rangle$ is an expression of the form $g=\sum_{i=1}^{k}p_{i}f_{i}q_{i}$, with $k\in\mathbb{N}$, $p_{i},q_{i}\in K\langle X\rangle$, and $f_{i}\in F$.
Such a representation serves as a certificate that $g\in(F)$.
Our library f4ncgb supports the computation of two kinds of cofactor representations for each output polynomial $g_{i}$: (i) w.r.t. the input $\{f_{1},\dots,f_{r}\}$ and all previously computed polynomials $\{g_{1},\dots,g_{i-1}\}$, or (ii) directly w.r.t. the input $\{f_{1},\dots,f_{r}\}$.
Both forms certify that the computed basis is indeed a subset of the input ideal (the latter directly, the former recursively). These representations can be derived during the row echelon reduction step of the F4 algorithm (step 1), when the transformation matrix $T$ is computed such that $T\cdot A=\text{rref}(A)$. Type (i) representations can be read off directly from the rows of $T$. Type (ii) requires an additional backward substitution step.
Remark 2
The backward substitution step for computing representations of type (ii) can be computationally expensive and lead to an exponential increase in the size of the representations. Thus, representations of type (i) are generally preferable. We nevertheless support both types as certain applications may require representations of type (ii).
8 Experiments
We evaluate f4ncgb against the related tools from Section 2, focusing on GBNP, Letterplace, Macaulay2, NCPoly, and operator_gb. We exclude bergman and Opal from the evaluation as they are no longer maintained and were already outperformed by earlier versions of Letterplace [30]. We also omit NCAlgebra as initial experiments indicated that it is several orders of magnitude slower than the other tools (NCAlgebra required almost 3 hours to solve braid3-11, while f4ncgb needed $0.1$ seconds). Due to licensing restrictions, we could not evaluate Magma. However, to offer a rough comparison, we include Magma timings from [33].
In Table 2, we list the runtimes of each tool on the benchmarks from [33], all of which are part of the SymbolicData project [39]. For f4ncgb, we use the default configuration: single-threaded execution with the tracer for multi-modular computations and the Mersenne prime optimization enabled, and proof logging disabled.
Table 3 reports timings for the parallelized version of f4ncgb. For these tests, we use (larger) instances from the SymbolicData project and we limit our comparison to Letterplace and Macaulay2, as all other tools performed significantly worse on the smaller examples. On these larger benchmarks, the single-threaded version of f4ncgb achieves speedups of up to $880\times$ over Letterplace and up to $510\times$ over Macaulay2. The parallelized version reaches speedups of up to $1500$\times$ and $920$\times$, respectively. All reported times are real (wall-clock) times.
All computations are performed over $\mathbb{Q}$ w.r.t. a degree-lexicographic monomial ordering. To ensure termination, we impose a degree bound on the considered ambiguities in all computations. The designated bounds are indicated by the number after the last “-” in the name of each example, e.g., in braid3-11 we only consider ambiguities of degree $\leq 11$.
All experiments were run with a timeout of $12\text{\,}\mathrm{h}$ (wall-clock time) on a cluster of dual-socket AMD EPYC 7313 @ $3.7\text{\,}\mathrm{GHz}$ machines running Ubuntu 24.04 with a memory limit of $30\,000\text{\,}\mathrm{MB}$ per computation.
Example f4ncgb Letterplace Macaulay2 Magma GBNP NCPoly operator_gb
lascala_neuh-10 $1.24$ $19.86$ $7.57$ $13.62$ $271.66$ $63.96$ $32.89$
serre-f4-15 $4.90$ $3.66$ $5.63$ $8.96$ $2262.04$ $4303.90$ $8404.25$
serre-ha11-15 $4.36$ $8.00$ $12.05$ $5.80$ $2348.60$ $6258.89$ $7793.97$
serre-eha112-13 $1.51$ $1.39$ $3.91$ $1.72$ $315.32$ $560.23$ $397.75$
4nilp5s-8 $3.24$ $24.48$ $18.91$ $9.54$ $1126.18$ $548.48$ $147.17$
braidXY-11 $15.86$ $87.97$ $1305.72$ $4.20$ $120.55$ $7649.53$ $5196.03$
ug2-x1x2x3x4-5 $0.18$ $2.56$ TO $0.83$ $2.61$ $26.33$ $6.53$
serre-e6-15 $21.39$ $14.80$ $14.51$ $40.99$ $2968.67$ $8862.33$ TO
braid3-11 $0.10$ $1.36$ $5.20$ $0.64$ $120.55$ $107.48$ $55.04$
ufn3-10 $2.60$ $34.76$ $6.05$ $2.26$ $244.30$ $284.16$ $58.67$
ls3nilp-10 $0.21$ $0.53$ $2.17$ $1.98$ $21.76$ $10.27$ $13.32$
Table 2: Timings (in sec) on benchmarks from [33]. For Magma, we report the timings from [33]. TO: $43\,200\text{\,}\mathrm{s}$\$
Example Letterplace Macaulay2 f4ncgb 1 core 4 cores 8 cores 16 cores 32 cores 4nilp5s-10 $1282.44$ $875.09$ $150.46$ $79.47$ $65.96$ $63.03$ $57.44$ braid3-16 18 953.1918\,953.19 14 291.3714\,291.37 $105.46$ $34.58$ $23.96$ $18.14$ $17.69$ braidX-18 TO TO $1977.27$ $601.12$ $364.29$ $260.65$ $239.30$ braidXY-12 $1847.97$ 18 887.6118\,887.61 $62.40$ $52.44$ $53.51$ $52.98$ $55.26$ holt_G3562h-17 TO TO 25 021.6225\,021.62 12 671.3812\,671.38 $8915.69$ $6824.36$ $5715.45$ lascala_neuh-13 $171.95$ $37.94$ $9.84$ $5.50$ $4.98$ $4.90$ $5.25$ lp1-15 24 166.8524\,166.85 33 923.1733\,923.17 $266.44$ $179.22$ $161.98$ $155.05$ $168.91$ lv2d10-100 TO 24 930.2524\,930.25 $48.43$ $27.68$ $26.98$ $47.65$ $85.93$ malle_G12h-100 $4142.27$ $163.38$ $89.93$ $74.15$ $74.19$ $73.91$ $79.86$ serre-e6-17 $113.95$ $84.99$ $130.62$ $130.53$ $131.01$ $128.97$ $133.14$ serre-ha11-17 $99.09$ $108.41$ $38.18$ $31.66$ $30.77$ $30.84$ $31.87$
Table 3: Timings on larger benchmarks (in sec). TO: $43\,200\text{\,}\mathrm{s}$\$
9 Conclusion and Future Work
We presented f4ncgb, a new open-source C++ library for Gröbner basis computations in free algebras and discussed its implementation details and design choices. Our experiments demonstrate that f4ncgb establishes a new state of the art for noncommutative Gröbner basis computations.
On some instances, however, f4ncgb seems to be rather memory inefficient. Reducing the memory footprint of the library is part of future work. We also plan to integrate signature-based Gröbner basis algorithms, which have recently been generalized to the noncommutative setting [24, 25] and have proven useful in applications [26]. Moreover, we want to extend the supported coefficient domains to include Euclidean domains, in particular, the integers.
Acknowledgements
We thank the anonymous reviewers for their feedback and comments. We are also grateful to Yulia Mukhina for bringing the implementation in Macaulay2 to our attention.